# Temporal Feature Matters: A Framework for Diffusion Model Quantization

### v3

@@ Abstract
The Diffusion models, widely used for image generation, face significant challenges related to their broad applicability due to prolonged inference times and high memory demands. Efficient Post-Training Quantization (PTQ) is crucial to address these issues. However, unlike traditional models, diffusion models critically rely on the time-step for the multi-round denoising. Typically, each time-step is encoded into a hypersensitive temporal feature by several modules. Despite this, existing PTQ methods do not optimize these modules individually. Instead, they employ unsuitable reconstruction objectives and complex calibration methods, leading to significant disturbances in the temporal feature and denoising trajectory, as well as reduced compression efficiency. To address these challenges, we introduce a novel quantization framework that includes three strategies: 1) TIB-based Maintenance: Based on our innovative Temporal Information Block (TIB) definition, Temporal Information-aware Reconstruction (TIAR) and Finite Set Calibration (FSC) are developed to efficiently align original temporal features. 2) Cache-based Maintenance: Instead of indirect and complex optimization for the related modules, pre-computing and caching quantized counterparts of temporal features are developed to minimize errors. 3) Disturbance-aware Selection: Employ temporal feature errors to guide a fine-grained selection between the two maintenance strategies for further disturbance reduction. This framework preserves most of the temporal information and ensures high-quality end-to-end generation. Extensive testing on various datasets, diffusion models and hardware confirms our superior performance and acceleration.
@@ 1 Introduction
Generative modeling is pivotal in machine learning, particularly in applications like image [29, 82, 22, 74, 23, 71], voice [80, 73], and text synthesis [5, 99]. Diffusion models have demonstrated remarkable proficiency in generating high-quality samples across diverse domains. In comparison to generative adversarial networks (GANs) [16] and variational autoencoders (VAEs) [39], diffusion models successfully sidestep issues such as model collapse and posterior collapse, offering a more stable training regimen. However, the substantial computational cost [58, 94] poses a critical bottleneck hampering the widespread adoption of diffusion models. Specifically, this cost mainly stems from two primary factors. First, these models typically require hundreds of denoising steps to generate images, rendering the procedure considerably slower than that of GANs. Prior efforts [82, 54, 40, 51] have tackled this challenge by seeking shorter and more efficient sampling trajectories, thereby reducing the number of necessary denoising steps. Second, the substantial parameters and complex architecture of diffusion models demand considerable time and memory resources for mobile device inference [7, 101], particularly for foundational models pre-trained on large-scale datasets, e.g., Stable Diffusion [74], SD-XL [71] and SD-XL-turbo [78]. Our work aims to address the latter challenge, focusing on the compression of diffusion models to enhance their efficiency and applicability.
@@
Quantization, a technique for mapping high-precision floating-point numbers to lower-precision counterparts, stands as the most prevalent method for model compression [61, 15, 62, 12, 4]. Among different quantization paradigms, post-training quantization (PTQ) [61, 89, 25] incurs lower overhead and is more user-friendly without the need for retraining or fine-tuning with a huge amount of training data. While PTQ on conventional models has undergone extensive study [61, 45, 89, 15], its adaptation to diffusion models has shown huge performance degradation, especially under low-bit settings. For instance, Q-Diffusion [44] exhibits 6.81 Fréchet Inception Distance (FID) [21] increasing on CelebA-HQ $256\times 256$ [32] under 4-bit weight and 8-bit activation quantization. We believe the reason they fail to achieve better results is that they all overlook the sampling data independence and uniqueness of hypersensitive temporal features, which are generated from time-step $t$ through a few modules, used to control the denoising trajectory in diffusion models. As a result, the temporal feature disturbance stemming from quantization significantly impacts model performance in the existing studies.
@@
To tackle temporal feature disturbance, we first find that the modules generating temporal features are independent of the sampling data and thus treat the whole modules as the Temporal Information Block (TIB). However, existing methods [19, 44, 79, 85, 81] do not separately optimize this block during the quantization process, causing temporal features to overfit to limited calibration data. Additionally, since the maximum time-step for denoising is a finite positive integer, the temporal feature and the activations during its generation form a finite set. Therefore, the optimal approach is to optimize each element in this set individually. Inspired by these observations and analyses, we propose two temporal feature maintenance techniques: 1) TIB-based Maintenance: Based on TIB, a novel quantization reconstruction approach, Temporal Information-aware Reconstruction (TIAR), is devised. It aims to reduce temporal feature error as the optimization objective while isolating the network’s components related to sampled data during weight adjustment. Besides, we also introduce a calibration strategy, Finite Set Calibration (FSC), for the finite set of the temporal feature and activation during its generation. 2) Cache-based Maintenance: Further leveraging the independent and finite nature of the temporal feature, we directly pre-compute these features, and then optimize and cache their quantized versions to reduce the disturbance. In detail, this novel design only requires reloading cached features, avoiding online generation during inference, which enhances latency in specific cases. Moreover, we employ our Disturbance-aware Selection taking advantage of both maintenance approaches to further eliminate temporal feature error. In addition, only one generation process before deployment can determine the selection results for every single temporal feature, respectively. This highly ensures the efficiency of our selection scenario. Finally, evaluating on multiple datasets, diverse tasks, advanced models, and different hardware, our novel framework for reducing temporal feature disturbance achieved nearly lossless model compression with high inference speed.
@@
  • We propose an advanced quantization framework, consisting of 1) TIB-based Maintenance: Temporal Information-aware Reconstruction (TIAR) and Finite Set Calibration (FSC). Both are based on a Temporal Information Block specially defined for diffusion models. 2) Cache-based Maintenance: Directly optimize and save quantized temporal feature, and then reload for inference. 3) Disturbance-aware Selection: fine-grained select our maintenance strategies offline employing temporal feature error.
@@
  • Extensive experiments demonstrate our superior performance. For example, our framework reduces the FID score by 5.61 under the w4a8 configuration for SD-XL [71]. Additionally, we deploy the quantized model on different hardware to demonstrate the inference acceleration, e.g., 2.20$\times$ and 5.76$\times$ speedup on Intel® Xeon® Gold 6448Y Processor and Nvidia H800 GPU for SD-XL, respectively.
@@
This paper extends our conference version [24] in several aspects. 1) We compare and fine-grainedly analyze the sensitivity and quantization-induced disturbance between temporal features and other non-temporal features. 2) We propose Cache-based Maintenance parallel to TIB-based Maintenance with high efficiency. It can achieve comparable performance and even lower inference costs in some cases. 3) We propose Disturbance-aware Selection fine-grainedly harness both maintenance strategies to further reduce disturbance in temporal features. 4) We apply our framework to advanced Text-to-Image (T2I) models, e.g., SD-XL [71], SD-XL-turbo [78], and FLUX.1-Schnell [42]. Moreover, we also apply our methods to the video generation model–OpenSora [103]. 5) We deploy our quantized model employing CUTLASS [33] with our customized kernel on Nvidia H800 GPU and OpenVino [26] on Intel® Xeon® Gold 6448Y Processor to show significant acceleration. Additionally, we also test the efficiency of the quantized model on Nvidia Jetson Nano Orin and iPhone 15 Pro Max. 6) We provide comprehensive experiments and more ablative studies to investigate the effectiveness of our methods.
2 Related Work
2.1 Efficient Diffusion Models
@@
Diffusion models have demonstrated the ability to generate high-quality images. However, their extensive iterative process coupled with the computational cost of denoising via networks has impeded wide-ranging applications. To tackle this challenge, prior research has explored efficient methods to expedite the denoising process [36, 90, 1, 88]. These methods fall into two categories: those requiring re-training and advanced samplers tailored for pre-trained models. The former category includes techniques like diffusion process learning [9, 57, 102, 14], noise scale adaptation [64, 38], knowledge distillation [77, 56, 35], and sample trajectory refinement [88, 43, 63]. While these methods can speed up sampling, re-training a diffusion model proves resource-intensive and time-consuming. In contrast, the latter category involves designing efficient samplers for pre-trained diffusion models, eliminating the need for re-training. Key methods in this category encompass analytical trajectory estimation [3, 2], implicit samplers [40, 82, 98], and differential equation (DE) solvers such as customized stochastic differential equations (SDE) [83, 28, 34] and ordinary differential equations (ODE) [54, 51, 97]. While these approaches can reduce sampling iterations, diffusion models’ extensive parameters, and computational complexity restrict their deployment in real-world settings. In this paper, our work focuses on diminishing the time and memory overhead of the single-step denoising process using low-bit quantization in a post-training manner, a method orthogonal to previous speedup techniques.
@@
Besides the above-mentioned studies, there are also some works, e.g., SnapFusion [46] and BitsFusion [84] aim to extremely compress the diffusion model with a mix of multiple techniques for deployment on edge devices. However, these methods also require extensive resources for training or fine-tuning.
2.2 Model Quantization
@@
Quantization is a predominant technique for minimizing storage and computational costs. It can be categorized into quantization-aware training (QAT) [15, 27, 53, 96, 105] and post-training quantization (PTQ) [25, 45, 49, 61, 89]. QAT requires intensive model training with substantial data and computational demands. Correspondingly, PTQ compresses models without re-training, making it a preferred method due to its minimal data requirements and easy deployment on real hardware. In PTQ, given a high-precision value $x$, we map it into discrete one $\hat{x}$ using uniform quantization expressed as:
@@
where $s$ is the quantization step size, $z$ is the zero offset, and $b$ is the target bit-width. The clamp function $\Phi(\cdot)$ clips the rounded value $\left\lfloor\frac{x}{s}\right\rceil+z$ within the range of $[0,2^{b}-1]$. However, naive quantization may lead to accuracy degradation, especially for low-bit quantization. Recent studies [45, 13, 50, 47] have explored innovative strategies based on reconstruction to preserve model performance after low-bit quantization. For instance, AdaRound [61] extends PTQ to 4-bit on traditional vision models using a novel rounding mechanism with layer-wise reconstruction [25, 87]. BRECQ [45] balances cross-layer dependency and generalization error by leveraging neural network blocks and employing block-wise reconstruction. Additionally, QDrop [89], considering the impact of activation quantization when reconstructing layer/block outputs.
@@
In contrast, the iterative denoising process in diffusion models presents new challenges for PTQ in comparison to traditional models. PTQ4DM [79] represents the initial attempt to quantize diffusion models to 8-bit, albeit with limited experiments and lower resolutions. Conversely, Q-Diffusion [44] achieves enhanced performance and is evaluated on a broader dataset range. Moreover, PTQD [19] eliminates quantization noise through correlated and residual noise correction. Notably, traditional single-time-step PTQ calibration methods are unsuitable for diffusion models due to significant activation distribution changes with each time-step [79, 44, 85, 81]. ADP-DM [85] proposes group-wise quantization across time-steps for diffusion models, and TDQ [81] introduces distinct quantization parameters for different time-steps. However, all of the above works overlook the specificity of hypersensitive temporal features. To address temporal feature disturbance in the aforementioned works, our study delves into the inducements of the phenomenon and introduces a novel PTQ framework for diffusion models, significantly enhancing quantized diffusion model performance.
3 Preliminaries
@@
Diffusion models. Diffusion models [82, 22] iteratively add Gaussian noise with a variance schedule $\beta_{1},\ldots,\beta_{T}\in(0,1)$ to data $\mathbf{x}_{0}\sim q(\mathbf{x})$ for $T$ times as sampling process, resulting in a sequence of noisy samples $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$. In DDPMs [22], the former sampling process is a Markov chain, taking the form:
$\begin{split}q(\mathbf{x}_{1:T}|\mathbf{x}_{0})&=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\\ q(\mathbf{x}_{t}|\mathbf{x}_{t-1})&=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),\end{split}$ (2)
@@
where $\alpha_{t}=1-\beta_{t}$. Conversely, the denoising process removes noise from a sample from Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to gradually generate high-fidelity images. Nevertheless, due to the unavailability of the true reverse conditional distribution $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, diffusion models approximate it via variational inference by learning a Gaussian distribution $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};{\boldsymbol{\mu}}_{\theta}(\mathbf{x}_{t},t),{\boldsymbol{\Sigma}}_{\theta}(\mathbf{x}_{t},t))$, the ${\boldsymbol{\mu}}_{\theta}$ can be derived by reparameterization trick as follows:
@@
where $\|\cdot\|^{2}_{F}$ denotes the Frobenius norm. The function $f_{i}(\cdot)$ represents the $i^{\text{th}}$ Residual Bottleneck Block, and $\widehat{f_{i}}(\cdot)$ denotes its quantized counterpart. Furthermore, in the following sections, we use $n$ to denote the total number of Residual Bottleneck Blocks in a single diffusion model. Unless otherwise stated, we adopt the aforementioned method as our baseline in Sec. 4.
@@
Temporal feature in diffusion models. Also shown in Fig. 1, time-step $t$ is encoded with time embed [^1] PyTorch time embed implementation in diffusion models. and then passes through the embedding layer [^2] PyTorch embedding layer implementation in diffusion models. in each Residual Bottleneck Block, resulting in a series of unique activations. In this paper, we denote these activations as temporal features. Notably, temporal features are independent of $\mathbf{x}_{t}$ and unrelated to other temporal features from different time-steps. For enhanced clarity in our notation: the time embed function is denoted as $h(\cdot)$, the embedding layer within the $i^{\text{th}}$ Residual Bottleneck Block as $g_{i}(\cdot)$, and the resultant $i^{\text{th}}$ temporal feature at time-step $t$ as $\mathbf{X}_{t,i}^{temporal}$. Moreover, as illustrated in Fig. 1, the relationship among these components is expressed as
$\mathbf{X}_{t,i}^{temporal}=g_{i}(h(t))$ (6)
Additionally, we have found that temporal features play a crucial role in the context of the diffusion model as they hold unique and substantial physical implications. These features contain temporal information that indicates the current image’s position along the denoising trajectory. Within the architecture of the UNet, each time-step is converted into these temporal features, which then guide the denoising process by applying them to the features of images generated in successive iterations.
4 Temporal Feature Matters
We organize this section as follows. Firstly, we observe the disturbance induced by the previous method of the highly sensitive temporal feature in Sec. 4.1. Further findings in Sec. 4.2 confirm the drastic impact. Concurrently, we analyze the inducements in Sec. 4.3. Finally, we propose our quantization framework in Sec. 4.4.
4.1 Disturbance Observations
Based on Sec. 3, we investigate how temporal features exhibit salient sensitivity of disturbance, and then we identify the phenomenon of temporal feature disturbance induced by quantization.
Temporal feature sensitivity. Beyond the special physical significance of temporal feature on image generation, we find these features are more sensitive than others (i.e., other activations in the model can endure greater disturbances with minimal impact on performance). To validate this, we first introduce non-temporal features $\mathbf{X}_{t,i}^{non-temperal}$ to represent activations except temporal features for every time-step $t$. Then, we randomly select $n$ non-temporal features  [^3] Hence, here $i=1,\ldots,n$, the same for the temporal feature. at $t$, and apply varying levels of random Gaussian noise to both (i.e., temporal and non-temporal features). The noise can be formulated as follows:
$\mathbf{\Delta}_{\lambda}=\lambda\mathbf{\Delta}$ (7)

### v4

@@ Abstract
Diffusion models, widely used for image generation, face significant challenges related to their broad applicability due to prolonged inference times and high memory demands. Efficient Post-Training Quantization (PTQ) is crucial to address these issues. However, unlike traditional models, diffusion models critically rely on the timestep for the multi-round denoising. Typically, each timestep is encoded into a hypersensitive temporal feature by several modules. Despite this, existing PTQ methods do not optimize these modules individually. Instead, they employ unsuitable reconstruction objectives and complex calibration methods, leading to significant disturbances in the temporal feature and denoising trajectory, as well as reduced compression efficiency. To address these challenges, we introduce a novel quantization framework that includes three strategies: 1) TIB-based Maintenance: Based on our innovative Temporal Information Block (TIB) definition, Temporal Information-aware Reconstruction (TIAR) and Finite Set Calibration (FSC) are developed to efficiently align original temporal features. 2) Cache-based Maintenance: Instead of indirect and complex optimization for the related modules, pre-computing and caching quantized counterparts of temporal features are developed to minimize errors. 3) Disturbance-aware Selection: Employ temporal feature errors to guide a fine-grained selection between the two maintenance strategies for further disturbance reduction. This framework preserves most of the temporal information and ensures high-quality end-to-end generation. Extensive testing on various datasets, diffusion models, and hardware confirms our superior performance and acceleration.
@@ I Introduction
Generative modeling is pivotal in machine learning, particularly in applications like image [31, 81, 22, 73, 23, 70], voice [79, 72], and text synthesis [4, 98]. Diffusion models have demonstrated remarkable proficiency in generating high-quality samples across diverse domains. In comparison to generative adversarial networks (GANs) [16] and variational autoencoders (VAEs) [40], diffusion models successfully sidestep issues such as model collapse and posterior collapse, offering a more stable training regimen. However, the substantial computational cost [57, 93] poses a critical bottleneck hampering the widespread adoption of diffusion models. Specifically, this cost mainly stems from two primary factors. First, these models typically require hundreds of denoising steps to generate images, rendering the procedure considerably slower than that of GANs. Prior efforts [81, 53, 41, 51] have tackled this challenge by seeking shorter and more efficient sampling trajectories, thereby reducing the number of necessary denoising steps. Second, the substantial parameters and complex architecture of diffusion models demand considerable time and memory resources for mobile device inference [6, 100], particularly for foundational models pre-trained on large-scale datasets, e.g., Stable Diffusion [73], SD-XL [70], and SD-XL-turbo [77]. Our work aims to address the latter challenge, focusing on the compression of diffusion models to enhance their efficiency and applicability.
@@
Quantization, a technique for mapping high-precision floating-point numbers to lower-precision counterparts, stands as the most prevalent method for model compression [60, 14, 61, 11, 3]. Among different quantization paradigms, post-training quantization (PTQ) [60, 89, 27] incurs lower overhead and is more user-friendly without the need for retraining or fine-tuning with a huge amount of training data. While PTQ on conventional models has undergone extensive study [60, 46, 89, 14], its adaptation to diffusion models has shown huge performance degradation, especially under low-bit settings. For instance, Q-Diffusion [45] exhibits $\mathbf{6.81}$ Fréchet Inception Distance (FID) [21] increasing on CelebA-HQ $256\times 256$ [34] under $4$-bit weight and $8$-bit activation quantization. We believe the reason they fail to achieve better results is that they all overlook the sampling data independence and uniqueness of hypersensitive temporal features, which are generated from timestep $t$ through a few modules, used to control the denoising trajectory in diffusion models. As a result, the temporal feature disturbance stemming from quantization significantly impacts model performance in the existing studies.
@@
To tackle temporal feature disturbance, we first find that the modules generating temporal features are independent of the sampling data and thus treat the whole module as the Temporal Information Block (TIB). However, existing methods [19, 45, 78, 84, 80] do not separately optimize this block during the quantization process, causing temporal features to overfit to limited calibration data. Additionally, since the maximum timestep for denoising is a finite positive integer, the temporal feature and the activations during its generation form a finite set. Therefore, the optimal approach is to optimize each element in this set individually. Inspired by these observations and analyses, we propose two temporal feature maintenance techniques: 1) TIB-based Maintenance: Based on TIB, a novel quantization reconstruction approach, Temporal Information-aware Reconstruction (TIAR), is devised. It aims to reduce temporal feature error as the optimization objective while isolating the network’s components related to sampled data during weight adjustment. Besides, we also introduce a calibration strategy, Finite Set Calibration (FSC), for the finite set of the temporal feature and activation during its generation. 2) Cache-based Maintenance: Further leveraging the independent and finite nature of the temporal feature, we directly pre-compute these features, and then optimize and cache their quantized versions to reduce the disturbance. In detail, this novel design only requires reloading cached features, avoiding online generation during inference, which enhances latency in specific cases. Moreover, we employ a 3) Disturbance-aware Selection, taking advantage of both maintenance approaches to further eliminate temporal feature error. In addition, only one generation process before deployment can determine the selection results for every single temporal feature, respectively. This highly ensures the efficiency of our selection scenario. Finally, evaluating on multiple datasets, diverse tasks, advanced models, and different hardware, our novel framework for reducing temporal feature disturbance achieved nearly lossless model compression with high inference speed.
@@
  • We propose an advanced quantization framework, consisting of 1) TIB-based Maintenance: Temporal Information-aware Reconstruction (TIAR) and Finite Set Calibration (FSC). Both are based on a Temporal Information Block specially defined for diffusion models. 2) Cache-based Maintenance: Directly optimize and save quantized temporal feature, and then reload for inference. 3) Disturbance-aware Selection: Fine-grainedly select our maintenance strategies offline employing temporal feature error.
@@
  • Extensive experiments demonstrate our superior performance. For example, our framework reduces the FID score by $\mathbf{5.61}$ under the W4444A8888 configuration for SD-XL [70]. Additionally, we deploy the quantized model on different hardware to demonstrate the inference acceleration, e.g., $\mathbf{2.20\times}$ and $\mathbf{5.76\times}$ speedup on Intel® Xeon® Gold $6448Y$ Processor and NVIDIA $H800$ GPU for SD-XL, respectively.
@@
This paper extends our conference version [24] in several aspects. 1) We compare and fine-grainedly analyze the sensitivity and quantization-induced disturbance between temporal and other non-temporal features. 2) We propose Cache-based Maintenance parallel to TIB-based Maintenance with high efficiency. It can achieve comparable performance and even lower inference costs in some cases. 3) We propose Disturbance-aware Selection, which fine-grainedly harnesses both maintenance strategies to further reduce disturbance in temporal features. 4) We apply our framework to advanced Text-to-Image (T2I) models, e.g., SD-XL [70], SD-XL-turbo [77], and FLUX.1-Schnell [43]. Moreover, we also apply our methods to the video generation model–OpenSora [102]. 5) We deploy our quantized model employing CUTLASS [35] with our customized kernel on NVIDIA $H800$ GPU and OpenVino [28] on Intel® Xeon® Gold $6448Y$ Processor to show significant acceleration. Additionally, we also test the efficiency of the quantized model on NVIDIA Jetson Nano Orin and iPhone $15$ Pro Max. 6) We provide comprehensive experiments and more ablative studies to investigate the effectiveness of our methods.
II Related Work
II-A Efficient Diffusion Models
@@
Diffusion models have demonstrated the ability to generate high-quality images. However, their extensive iterative process, coupled with the computational cost of denoising via networks, has impeded wide-ranging applications. To tackle this challenge, prior research has explored efficient methods to expedite the denoising process [88, 26, 87]. These methods fall into two categories: those requiring re-training and advanced samplers tailored for pre-trained models. The former category includes techniques like diffusion process learning [8, 56, 101, 13], noise scale adaptation [63, 39], knowledge distillation [76, 55, 37], and sample trajectory refinement [88, 44, 62]. While these methods can speed up sampling, re-training a diffusion model proves resource-intensive and time-consuming. In contrast, the latter category involves designing efficient samplers for pre-trained diffusion models, eliminating the need for re-training. Key methods in this category encompass analytical trajectory estimation [2, 1], implicit samplers [41, 81, 97], and differential equation (DE) solvers such as customized stochastic differential equations (SDE) [82, 30, 36] and ordinary differential equations (ODE) [53, 51, 96]. While these approaches can reduce sampling iterations, diffusion models’ extensive parameters, and computational complexity restrict their deployment in real-world settings. In this paper, our work focuses on diminishing the time and memory overhead of the single-step denoising process using low-bit quantization in a post-training manner, a method orthogonal to previous speedup techniques.
@@
Besides the above-mentioned studies, there are also some works, e.g., SnapFusion [47] and BitsFusion [83] aim to extremely compress the diffusion model with a mix of multiple techniques for deployment on edge devices. However, these methods also require extensive training resources.
II-B Model Quantization
@@
Quantization is a predominant technique for minimizing storage and computational costs. It can be categorized into quantization-aware training (QAT) [14, 95, 104, 25] and post-training quantization (PTQ) [46, 60, 89, 15]. QAT requires intensive model training with substantial data and computational demands. Correspondingly, PTQ compresses models without re-training, making it a preferred method due to its minimal data requirements and easy deployment on real hardware. In PTQ, given a high-precision value $x$, we map it into discrete one $\hat{x}$ using uniform quantization expressed as:
@@
where $s$ is the quantization step size, $z$ is the zero offset, and $b$ is the target bit-width. The clamp function $\Phi(\cdot)$ clips the rounded value $\left\lfloor\frac{x}{s}\right\rceil+z$ within the range of $[0,2^{b}-1]$. However, naive quantization may lead to accuracy degradation, especially for low-bit quantization. Recent studies [46, 12, 50, 48] have explored innovative strategies based on reconstruction to preserve model performance after low-bit quantization. For instance, AdaRound [60] extends PTQ to $4$-bit on traditional vision models using a novel rounding mechanism with layer-wise reconstruction [27, 86]. BRECQ [46] balances cross-layer dependency and generalization error by leveraging neural network blocks and employing block-wise reconstruction. Additionally, QDrop [89] considers the impact of activation quantization during reconstruction.
@@
In contrast, the iterative denoising process in diffusion models presents new challenges for PTQ in comparison to traditional models. PTQ4DM [78] represents the initial attempt to quantize diffusion models to $8$-bit, albeit with limited experiments and lower resolutions. Conversely, Q-Diffusion [45] achieves enhanced performance and is evaluated on a broader dataset range. Moreover, PTQD [19] eliminates quantization noise through correlated and residual noise correction. Notably, traditional single-timestep PTQ calibration methods are unsuitable for diffusion models due to significant activation distribution changes with each timestep [78, 45, 84, 80]. ADP-DM [84] proposes group-wise quantization across timesteps for diffusion models, and TDQ [80] introduces distinct quantization parameters for different timesteps. However, all of the above works overlook the specificity of hypersensitive temporal features. To address temporal feature disturbance in the aforementioned works, our study delves into the inducements of the phenomenon and introduces a novel PTQ framework for diffusion models, significantly enhancing quantized diffusion model performance.
III Preliminaries
@@
Diffusion models. Diffusion models [81, 22] iteratively add Gaussian noise with a variance schedule $\beta_{1},\ldots,\beta_{T}\in(0,1)$ to data $\mathbf{x}_{0}\sim q(\mathbf{x})$ for $T$ times as sampling process, resulting in a sequence of noisy samples $\mathbf{x}_{1},\ldots,\mathbf{x}_{T}$. In DDPMs [22], the sampling process is a Markov chain, taking the form:
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$ $=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$ (2)
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ $=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$ (3)
@@
where $\alpha_{t}=1-\beta_{t}$. Conversely, the denoising process removes noise from a sample of Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to gradually generate high-fidelity images. Nevertheless, due to the unavailability of the true reverse conditional distribution $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, diffusion models approximate it via variational inference by learning a Gaussian distribution $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};{\boldsymbol{\mu}}_{\theta}(\mathbf{x}_{t},t),{\boldsymbol{\Sigma}}_{\theta}(\mathbf{x}_{t},t))$, the ${\boldsymbol{\mu}}_{\theta}$ can be derived by reparameterization trick as follows:
@@
where $\|\cdot\|^{2}_{F}$ denotes the Frobenius norm. The function $f_{i}(\cdot)$ represents the $i$-th Residual Bottleneck Block, and $\widehat{f_{i}}(\cdot)$ denotes its quantized counterpart. Furthermore, in the following sections, we use $n$ to denote the total number of Residual Bottleneck Blocks in a single diffusion model. Unless otherwise stated, we adopt the aforementioned method as our baseline in Sec. IV.
@@
Temporal feature in diffusion models. Also shown in Fig. 1, timestep $t$ is encoded with time embed [^1] PyTorch time embed implementation in diffusion models. and then passes through the embedding layer [^2] PyTorch embedding layer implementation in diffusion models. in each Residual Bottleneck Block, resulting in a series of unique activations. In this paper, we denote these activations as temporal features. Notably, temporal features are independent of $\mathbf{x}_{t}$ and unrelated to other temporal features from different timesteps. For enhanced clarity in our notation: the time embed function is denoted as $h(\cdot)$, the embedding layer within the $i$-th Residual Bottleneck Block as $g_{i}(\cdot)$, and the resultant $i$-th temporal feature at timestep $t$ as $\mathbf{X}_{t,i}^{temporal}$. As illustrated in Fig. 1, the relationship among these components is expressed as
$\mathbf{X}_{t,i}^{temporal}=g_{i}(h(t)).$ (7)
Additionally, we have found that temporal features play a crucial role in the context of the diffusion model as they hold unique and substantial physical implications. These features contain temporal information that indicates the current image’s position along the denoising trajectory. Within the architecture of the UNet, each timestep is converted into these temporal features, which then guide the denoising process by applying them to the image features generated in successive iterations.
IV Temporal Feature Matters
We organize this section as follows. Firstly, we observe the disturbance induced by the previous method of the highly sensitive temporal feature in Sec. IV-A. Further findings in Sec. IV-B confirm the drastic impact. Concurrently, we analyze the inducements in Sec. IV-C. Finally, we propose our quantization framework in Sec. IV-D.
IV-A Disturbance Observations
Based on Sec. III, we investigate how temporal features exhibit salient sensitivity to disturbance, and then we identify the phenomenon of temporal feature disturbance induced by quantization.
Temporal feature sensitivity. Beyond the special physical significance of temporal feature on image generation, we find these features are more sensitive than others (i.e., other activations in the model can endure greater disturbances with minimal impact on performance). To validate this, we first introduce non-temporal features $\mathbf{X}_{t,i}^{non-temporal}$ to represent activations except temporal features for every timestep $t$. Then, we randomly select $n$ non-temporal features  [^3] Hence, here $i=1,\ldots,n$, the same for the temporal feature. at $t$, and apply varying levels of random Gaussian noise to both (i.e., temporal and non-temporal features). The noise can be formulated as follows:
$\mathbf{\Delta}_{\lambda}=\lambda\mathbf{\Delta},$ (8)